OpenStudy (anonymous):

What degree Fahrenheit is it if the temperature is 15 degrees Celsius? Show all steps or reasoning. Be sure to check your answer for reasonableness. What degree Fahrenheit is it if the temperature is 15 degrees Celsius? Show all steps or reasoning. Be sure to check your answer for reasonableness. @Mathematics

OpenStudy (jamesj):

Celsius is a scale based on 0 = freezing temp of water 100 = boiling temp of water Fahrenheit has of course 32 and 212, a difference of 180. Hence to get from temp in celsius, call it C, to to temp in fahrenheit, call it F, we need to scale the temperature by 180/100 and then add 32. I.e., F = 180/100C + 32 = 9/5 . C + 32 Therefore in your case, C = 15. Now substitute and find the temp F.

OpenStudy (jamesj):

You can check for reasonableness by making sure that your answer is between 32 and 212, which it had better be, and somewhat closer to 32 than to 212.

OpenStudy (anonymous):

can you help me with the other one i posted??